# Sanjay Pradeepa, Chen Wangb, Matthew M. Dahmc, Jeff D. Eldredgea, and Candace S. J. Tsaid

### v1

@@ 1 Introduction
Carbon nanotubes (CNTs) and carbon nanofibers (CNFs) are a rapidly expanding class of engineered carbon structures whose exceptional mechanical, electrical, and thermal properties are being harnessed across diverse technologies34. From carbon-fiber bicycle frames and sporting goods to advanced electronics, aerospace structures, and energy devices, CNTs and related carbon nanomaterials (CNMs) are increasingly embedded in products across both consumer and high-technology sectors. As these materials transition from laboratory-scale research to large-volume industrial production and widespread use, the potential for human and environmental exposure increases accordingly. This is a significant concern, as toxicological studies have indicated that certain types of CNTs, particularly those with high aspect ratios, can elicit harmful biological responses upon inhalation, including persistent pulmonary inflammation and fibrosis, drawing parallels to asbestos-like pathogenicity 12, 10. Unlike many other engineered nanomaterials, carbon nanomaterials can differ greatly in their physical and chemical properties depending on how they are synthesized and treated after production, and recent epidemiological and toxicological studies suggest that fiber dimensions and the surface properties of agglomerates or bundles are important determinants of their toxic effects 22, 6, 24. These findings underscore the critical need for improved particle identification, counting, and classification protocols to enable meaningful exposure and risk assessment. Accurate morphological characterization of airborne particles is essential for occupational health surveillance, particularly for quantifying the respirable fraction and evaluating the biological hazard potential of CNMs in workplace environments where exposure may occur.
@@
Scanning Electron Microscopy (SEM) and Transmission Electron Microscopy (TEM) are the typical visualization methods for characterizing micron to nanometer sized carbon nanomaterials, particularly in occupational exposure studies where particles are collected on sampling media such as filters, mesh or TEM grids according to established sampling methods or NIOSH protocols. While these techniques yield images rich in quantitative and qualitative data, extracting actionable insights presents a significant and well-documented bottleneck. The analysis—comprising particle identification, quantification, and morphological classification—traditionally relies on manual or semi-automated workflows using tools like ImageJ 9 and Fiji30. To isolate particles from porous filters or complex backgrounds, researchers must often develop complex, multi-step image processing workflows involving advanced thresholding, morphological operations and watershedding. These workflows are notoriously brittle, often failing when applied to images with varying illumination, background textures, or particle agglomeration states. Consequently, this process is not only exceptionally time-consuming but is also subject to significant inter-operator variability, limiting the throughput and statistical reliability required for robust risk assessment.
@@
To overcome the limitations of manual analysis, the field has increasingly turned to machine learning (ML) and computer vision. Early efforts demonstrated the potential of Convolutional Neural Networks (CNNs) for these tasks36, 19, 8, 13, 46, 44, 37. For instance, deep learning based detection and segmentation models such as YOLO 27 and U-Net29 have been adapted for particle detection and pixel-wise segmentation in SEM and TEM images33, 31, but often struggle with the dense, overlapping, and morphologically diverse nature of nanoparticle agglomerates, and typically require extensive, manually-annotated bounding boxes for training28, 45. For the more complex task of classification, Luo et al. (2021) 20 introduced a transfer learning approach that combined a VGG-16 architecture 35 with a Vector of Locally Aggregated Descriptors (VLAD) 2 to classify TEM images of CNTs. While a significant step toward automation, this method highlighted a key challenge: its feature extraction, which sampled random pixels from the entire image, required a large dataset of over 5300 images and extensive data augmentation to achieve high accuracy, and cannot be applied to images with multiple types of CNT. The high data dependency of these models has remained a significant barrier to the accessible deployment of deep learning models in materials science and occupational exposure labs.
@@
The recent advent of large-scale, pre-trained foundation models for computer vision, however, represents a paradigm shift. Models like the Segment Anything Model (SAM) 17 are trained on millions of images with billions of segmentation masks, enabling them to perform "zero-shot" or "few-shot" segmentation of any object in an image with remarkable precision. This technology has already been recognized as a powerful tool for scientific imaging. Archit et al. (2025) 3 demonstrated this by developing "Segment Anything for Microscopy (µSAM)," a fine-tuned version of SAM for light and electron microscopy that can be used as an interactive tool, effectively replacing the laborious manual segmentation of traditional software. Concurrently, self-supervised models like DINOv2 25 have learned to produce highly discriminative feature representations from images without human labels. This combination of segmentation and feature extraction using foundation models has the potential to be a state-of-the-art approach. For example, Barnatan et al. (2025) 4 recently proposed a zero-shot shape classification pipeline for nanoparticles by combining SAM-generated segments with DINOv2 feature embeddings, further validating the power of this coupled approach.
@@
In this work, we harness these synergistic foundation models to build a comprehensive framework for both the quantification and classification of carbon nanotubes from electron micrographs. Our approach is a multi-stage process designed to first enable rapid, high-quality segmentation and quantification, and then to leverage these segments for robust, data-efficient classification. Particle segmentation is already a necessary step in the quantification workflow for counting and measuring individual CNT structures; our methodology integrates automated morphological classification directly into this existing process, allowing researchers to obtain both quantitative metrics and categorical labels simultaneously without additional manual effort. We present and validate a semi-automated, SAM-based interactive tool for the rapid and accurate segmentation of CNTs from diverse SEM and TEM images as shown in Fig 1. This tool is designed to replace traditional manual methods, allowing researchers with no computational background to perform robust particle counting and sizing. We then introduce a novel, data-efficient classification pipeline that leverages DINOv2 (ViT-B/14) for feature extraction. By guiding the DINOv2 encoder to extract features only from the segmented particles, we eliminate background noise and enable a simple MLP classifier to achieve state-of-the-art performance in categorizing particles into classes such as fiber, cluster, matrix, and matrix surface as shown in Fig 2. We demonstrate that this segmentation-guided pipeline achieves 95.5% classification accuracy while using only a fraction of the training data required by previous benchmarks 20, offering a generalizable and accessible framework for the rapid and reliable quantification and characterization of CNTs.
@@ 2 Dataset and Methods
This work uses two distinct, purpose-built datasets designed to validate our segmentation framework and train our classification pipeline. For assessing the segmentation performance of the interactive SAM tool, we curated an in-house collection of 200 electron microscope images evenly distributed between Scanning Electron Microscopy (SEM) and Transmission Electron Microscopy (TEM) modalities. This segmentation validation dataset deliberately captures the full spectrum of imaging conditions and sample characteristics encountered in CNT exposure assessment 22, including samples collected on polycarbonate filters from open-face and closed-face cassettes 23, 5, and Tsai Diffusion Samplers 38, as well as TEM grids placed on the filter. The images exhibit substantial morphological diversity from isolated fibers to complex clusters and dense agglomerates, alongside considerable variation in substrate texture and background contrast, as seen in Fig. 1, both of which commonly challenge conventional segmentation algorithms. For each image, we generated detailed ground-truth segmentation masks through a combination of Fiji-based segmentation and manual refinement to precisely delineate particle boundaries.
For the classification task, we employed a subset of 1,800 images from a comprehensive repository of 5,300 grayscale TEM images of various carbon nanomaterials collected by the U.S. National Institute for Occupational Safety and Health (NIOSH) 43. We generated segmentation masks for each selected image using the interactive SAM tool and assigned expert-annotated labels corresponding to one of four morphologically distinct classes that reflect the native structures of CNTs/CNFs and alternative structural forms commonly observed in field samples. A "fiber" represents cylindrical carbon nanostructures with various stacking arrangements of graphene sheets and aspect ratios exceeding 3:1, consistent with NIOSH counting criteria for airborne fibers associated with elevated toxicological potential 22. A "cluster" denotes bundles of such fibers that are loosely contacted, formed through inter-fiber interactions and the inherent flexibility of CNT/CNFs, which are known to agglomerate easily. A "matrix" describes the most common morphology in workplace air samples, where CNT/CNFs agglomerate and pack into condensed, web-like structures that obscure individual particle boundaries. Finally, a "matrix surface" identifies CNT/CNF fibers embedded within large particles comprising other carbonaceous materials, catalytic particles, and intermediate products emitted during manufacturing operations, where tube/fiber structures extrude from the particle surface and remain discernible at higher magnifications.
@@ 2.1 Segmentation Pipeline Architecture
In the first step, the SAM-based interactive tool implements a streamlined click-based workflow where users load electron microscope images. When scale bars are present (as is standard in exposure assessment studies), the system automatically extracts calibration information using easyOCR 16 text detection, eliminating manual scale entry. Upon loading, the image is passed through the SAM ViT (Vision Transformer) encoder to generate a high-dimensional feature embedding, a computationally intensive operation performed only once per image and cached for all subsequent interactions. Users then interact with the image through intuitive prompts: positive clicks (left-click) on target particles and optional negative clicks (right-click) to exclude undesired regions. With each prompt, the lightweight decoder operates on the pre-computed embedding to generate segmentation masks in milliseconds, enabling real-time iterative refinement until precise particle boundaries are achieved. The resulting masks undergo automated post-processing using scikit-image 42 morphological operations including remove_small_objects() for noise filtering and custom functions to eliminate edge artifacts. Connected component analysis via measure.label() identifies individual particles, and regionprops() extracts comprehensive morphological descriptors for each segmented object. Our approach leverages SAM’s powerful zero-shot capabilities within a human-in-the-loop workflow, removing the need for per-image parameter tuning while remaining accessible to users without scientific image analysis expertise.
@@ 3.2 Classification Performance
We systematically evaluated 24 model configurations combining two foundation models (SAM ViT-B/16 and DINOv2 ViT-B/14), two sampling strategies (mask-guided and uniform), three pooling methods (average, max, and avg+max concatenation), and two classifier types (linear and MLP). Test accuracy across all configurations ranged from 86.6% to 95.5%, with a mean of 91.4% $\pm$ 2.4% (Table 2; complete results in ESI Table S1). The best-performing architecture—DINOv2 with mask-guided sampling, avg+max pooling, and MLP classifier—achieved 95.5% test accuracy, substantially outperforming the baseline method of Luo et al.20 (90.9%) despite using only 34% of their training data (1,620 vs 4,800 images). This improvement demonstrates that foundation models pre-trained on large-scale vision datasets can effectively transfer to specialized nanomaterial classification tasks with relatively small labeled datasets, and that segmentation-guided feature extraction provides benefits over uniform sampling approaches.
@@
To understand why DINOv2 substantially outperforms both traditional CNNs and SAM for CNT classification, we visualized the hierarchical feature representations learned by the model across different transformer layers (Figure 7). These activation maps reveal how DINOv2 progressively builds increasingly abstract and task-specific representations of CNT morphologies 26, 15. Layer 1 activations respond primarily to low-level visual features—edges, texture gradients, and local contrast variations—producing bright activation along fiber boundaries, particle edges, and regions of high structural detail. These early-layer features are largely task-agnostic, capturing primitive visual elements present in any natural or microscopy image. By Layer 6, these distinct spatial boundaries begin to diffuse as the transformer integrates information across image patches; the resulting grid-like patterns indicate that the model is mixing local textures into distributed mid-level representations rather than encoding explicit object boundaries at this stage. The most striking transformation occurs in Layer 11, where we shift from visualizing raw activations to visualizing CLS token similarity - a proxy for the model’s "visual attention". These maps show that the model recovers spatial specificity but with semantic meaning: it activates strongly and exclusively on CNT particles while actively suppressing background regions, and imaging artifacts like TEM grid bars. This selective focus emerges naturally from self-supervised pre-training, confirming the emergent segmentation properties previously observed in DINO-based architectures7, 11, which effectively learn to "look" only at the relevant nanomaterial even without explicit segmentation training.
@@
To visualize how the MLP classifier transforms DINOv2’s foundation model features into highly discriminative representations, we applied t-SNE 39, 40, 41 dimensionality reduction to embeddings at three stages of the classification pipeline (Figure 8). The raw 3840-dimensional DINOv2 hypercolumn features (Figure 8a) already exhibit meaningful structure: the four CNT morphology classes form loosely separated clusters, with Fiber and Matrix samples occupying relatively distinct regions while Fiber and Cluster samples show greater overlap, consistent with their morphological similarity in a few images. This initial separation validates that DINOv2’s self-supervised pre-training on natural images transfers effectively to nanomaterial microscopy, capturing morphology-relevant features without task-specific fine-tuning. However, substantial inter-class overlap remains, particularly between Fiber and Cluster samples, and between Matrix and MatrixSurface samples explaining the 2% classification error in the confusion matrix. The first MLP layer (512 dimensions, Figure 8b) significantly refines this structure: class clusters become more compact and better separated, with decision boundaries beginning to form between previously overlapping regions. The most dramatic transformation occurs after the second MLP layer (128 dimensions, 8c), where all four classes form tight, well-separated manifolds with minimal overlap. This progression demonstrates that the MLP learns to amplify task-relevant dimensions of the DINOv2 feature space while suppressing irrelevant variations, effectively discovering a low-dimensional embedding optimized for CNT morphology discrimination. The small number of outlier points visible in each cluster correspond precisely to the misclassified samples in the confusion matrix, representing genuinely ambiguous borderline cases rather than random errors. Notably, even in the final 128-dimensional space, MatrixSurface and Cluster outliers tend to drift toward each other’s clusters, while Fiber and Matrix remain distinctly separated, confirming that the 2% Cluster-to-MatrixSurface confusion reflects true morphological ambiguity along a continuum rather than model failure. This visualization provides interpretable evidence that our model’s 95.5% test accuracy stems from learning a geometrically meaningful feature space where morphological similarity corresponds directly to proximity in the embedding, rather than memorizing superficial image patterns.

### v2

@@ 1 Introduction
Carbon nanotubes (CNTs) and carbon nanofibers (CNFs) are a rapidly expanding class of engineered carbon structures whose exceptional mechanical, electrical, and thermal properties are being harnessed across diverse technologies38. From carbon-fiber bicycle frames and sporting goods to advanced electronics, aerospace structures, and energy devices, CNTs and related carbon nanomaterials (CNMs) are increasingly embedded in products across both consumer and high-technology sectors. As these materials transition from laboratory-scale research to large-volume industrial production and widespread use, the potential for human and environmental exposure increases accordingly. This is a significant concern, as toxicological studies have indicated that certain types of CNTs, particularly those with high aspect ratios, can elicit harmful biological responses upon inhalation, including persistent pulmonary inflammation and fibrosis, drawing parallels to asbestos-like pathogenicity 15, 13. Unlike many other engineered nanomaterials, carbon nanomaterials can differ greatly in their physical and chemical properties depending on how they are synthesized and treated after production, and recent epidemiological and toxicological studies suggest that fiber dimensions and the surface properties of agglomerates or bundles are important determinants of their toxic effects 34, 25, 7, 27. These findings underscore the critical need for improved particle identification, counting, and classification protocols to enable meaningful exposure and risk assessment. Accurate morphological characterization of airborne particles is essential for occupational health surveillance, particularly for quantifying the respirable fraction and evaluating the biological hazard potential of CNMs in workplace environments where exposure may occur.
@@
Scanning Electron Microscopy (SEM) and Transmission Electron Microscopy (TEM) are the typical visualization methods for characterizing micron to nanometer sized carbon nanomaterials, particularly in occupational exposure studies where particles are collected on sampling media such as filters, mesh or TEM grids according to established sampling methods or NIOSH protocols7. While these techniques yield images rich in quantitative and qualitative data, extracting actionable insights presents a significant and well-documented bottleneck. The analysis—comprising particle identification, quantification, and morphological classification—traditionally relies on manual or semi-automated workflows using tools like ImageJ 10 and Fiji33. To isolate particles from porous filters or complex backgrounds, researchers must often develop complex, multi-step image processing workflows involving advanced thresholding, morphological operations and watershedding. These workflows are notoriously brittle, often failing when applied to images with varying illumination, background textures, or particle agglomeration states. Consequently, this process is not only exceptionally time-consuming but is also subject to significant inter-operator variability, limiting the throughput and statistical reliability required for robust risk assessment.
@@
To overcome the limitations of manual analysis, the field has increasingly turned to machine learning (ML) and computer vision. Early efforts demonstrated the potential of Convolutional Neural Networks (CNNs) for these tasks40, 22, 9, 16, 50, 48, 41. For instance, deep learning based detection and segmentation models such as YOLO 30 and U-Net32 have been adapted for particle detection and pixel-wise segmentation in SEM and TEM images37, 35, but often struggle with the dense, overlapping, and morphologically diverse nature of nanoparticle agglomerates, and typically require extensive, manually-annotated bounding boxes for training31, 49. For the more complex task of classification, Luo et al. (2021) 23 introduced a transfer learning approach that combined a VGG-16 architecture 39 with a Vector of Locally Aggregated Descriptors (VLAD) 2 to classify TEM images of CNTs. While a significant step toward automation, this method highlighted a key challenge: its feature extraction, which sampled random pixels from the entire image, required a large dataset of over 5300 images and extensive data augmentation to achieve high accuracy, and cannot be applied to images with multiple types of CNT. The high data dependency of these models has remained a significant barrier to the accessible deployment of deep learning models in materials science and occupational exposure labs.
@@
The recent advent of large-scale, pre-trained foundation models for computer vision, however, represents a paradigm shift. Models like the Segment Anything Model (SAM) 20 are trained on millions of images with billions of segmentation masks, enabling them to perform "zero-shot" or "few-shot" segmentation of any object in an image with remarkable precision. This technology has already been recognized as a powerful tool for scientific imaging. Archit et al. (2025) 3 demonstrated this by developing "Segment Anything for Microscopy (µSAM)," a fine-tuned version of SAM for light and electron microscopy that can be used as an interactive tool, effectively replacing the laborious manual segmentation of traditional software. Concurrently, self-supervised models like DINOv2 28 have learned to produce highly discriminative feature representations from images without human labels. This combination of segmentation and feature extraction using foundation models has the potential to be a state-of-the-art approach. For example, Barnatan et al. (2025) 4 recently proposed a zero-shot shape classification pipeline for nanoparticles by combining SAM-generated segments with DINOv2 feature embeddings, further validating the power of this coupled approach.
@@
In this work, we harness these synergistic foundation models to build a comprehensive framework for both the quantification and classification of carbon nanotubes from electron micrographs. Our approach is a multi-stage process designed to first enable rapid, high-quality segmentation and quantification, and then to leverage these segments for robust, data-efficient classification. Particle segmentation is already a necessary step in the quantification workflow for counting and measuring individual CNT structures; our methodology integrates automated morphological classification directly into this existing process, allowing researchers to obtain both quantitative metrics and categorical labels simultaneously without additional manual effort. We present and validate a semi-automated, SAM-based interactive tool for the rapid and accurate segmentation of CNTs from diverse SEM and TEM images as shown in Fig 1. This tool is designed to replace traditional manual methods, allowing researchers with no computational background to perform robust particle counting and sizing. We then introduce a novel, data-efficient classification pipeline that leverages DINOv2 (ViT-B/14) for feature extraction. By guiding the DINOv2 encoder to extract features only from the segmented particles, we eliminate background noise and enable a simple MLP classifier to achieve state-of-the-art performance in categorizing particles into classes such as fiber, cluster, matrix, and matrix surface as shown in Fig 2. We demonstrate that this segmentation-guided pipeline achieves 95.5% classification accuracy while using only a fraction of the training data required by previous benchmarks 23, offering a generalizable and accessible framework for the rapid and reliable quantification and characterization of CNTs.
@@ 2 Dataset and Methods
[…]
For the classification task, we employed a subset of 1,800 images from a comprehensive repository of 5,300 grayscale TEM images of various CNT/CNF materials from 26 facilities collected by the U.S. National Institute for Occupational Safety and Health (NIOSH) 47, 12, 11. By utilizing this dataset, our study targets the classification of complex industrial field samples, as opposed to lab-generated samples with distinct morphological features. We generated segmentation masks for each selected image using the interactive SAM tool and assigned expert-annotated labels corresponding to one of four morphologically distinct classes that reflect the native structures of CNTs/CNFs and alternative structural forms commonly observed in field samples. A "fiber" represents cylindrical carbon nanostructures with various stacking arrangements of graphene sheets and aspect ratios exceeding 3:1, consistent with NIOSH counting criteria for airborne fibers associated with elevated toxicological potential 5. A "cluster" denotes bundles of such fibers that are loosely contacted, formed through inter-fiber interactions and the inherent flexibility of CNT/CNFs, which are known to agglomerate easily. A "matrix" describes the most common morphology in workplace air samples, where CNT/CNFs agglomerate and pack into condensed, web-like structures that obscure individual particle boundaries. Finally, a "matrix surface" identifies CNT/CNF fibers embedded within large particles comprising other carbonaceous materials, catalytic particles, and intermediate products emitted during manufacturing operations, where tube/fiber structures extrude from the particle surface and remain discernible at higher magnifications7.
@@ 2.1 Segmentation Pipeline Architecture
In the first step, the SAM-based interactive tool implements a streamlined click-based workflow where users load electron microscope images. When scale bars are present (as is standard in exposure assessment studies), the system automatically extracts calibration information using easyOCR 19 text detection, eliminating manual scale entry. Upon loading, the image is passed through the SAM ViT (Vision Transformer) encoder to generate a high-dimensional feature embedding, a computationally intensive operation performed only once per image and cached for all subsequent interactions. Users then interact with the image through intuitive prompts: positive clicks (left-click) on target particles and optional negative clicks (right-click) to exclude undesired regions. With each prompt, the lightweight decoder operates on the pre-computed embedding to generate segmentation masks in milliseconds, enabling real-time iterative refinement until precise particle boundaries are achieved. The resulting masks undergo automated post-processing using scikit-image 46 morphological operations including remove_small_objects() for noise filtering and custom functions to eliminate edge artifacts. Connected component analysis via measure.label() identifies individual particles, and regionprops() extracts comprehensive morphological descriptors for each segmented object. Our approach leverages SAM’s powerful zero-shot capabilities within a human-in-the-loop workflow, removing the need for per-image parameter tuning while remaining accessible to users without scientific image analysis expertise.
@@ 3.2 Classification Performance
We systematically evaluated 24 model configurations combining two foundation models (SAM ViT-B/16 and DINOv2 ViT-B/14), two sampling strategies (mask-guided and uniform), three pooling methods (average, max, and avg+max concatenation), and two classifier types (linear and MLP). Test accuracy across all configurations ranged from 86.6% to 95.5%, with a mean of 91.4% $\pm$ 2.4% (Table 2; complete results in ESI Table 5). The best-performing architecture—DINOv2 with mask-guided sampling, avg+max pooling, and MLP classifier—achieved 95.5% test accuracy, substantially outperforming the baseline method of Luo et al.23 (90.9%) despite using only 34% of their training data (1,620 vs 4,800 images). This improvement demonstrates that foundation models pre-trained on large-scale vision datasets can effectively transfer to specialized nanomaterial classification tasks with relatively small labeled datasets, and that segmentation-guided feature extraction provides benefits over uniform sampling approaches.
@@
To understand why DINOv2 substantially outperforms both traditional CNNs and SAM for CNT classification, we visualized the hierarchical feature representations learned by the model across different transformer layers (Figure 7). These activation maps reveal how DINOv2 progressively builds increasingly abstract and task-specific representations of CNT morphologies 29, 18. Layer 1 activations respond primarily to low-level visual features—edges, texture gradients, and local contrast variations—producing bright activation along fiber boundaries, particle edges, and regions of high structural detail. These early-layer features are largely task-agnostic, capturing primitive visual elements present in any natural or microscopy image. By Layer 6, these distinct spatial boundaries begin to diffuse as the transformer integrates information across image patches; the resulting grid-like patterns indicate that the model is mixing local textures into distributed mid-level representations rather than encoding explicit object boundaries at this stage. The most striking transformation occurs in Layer 11, where we shift from visualizing raw activations to visualizing CLS token similarity - a proxy for the model’s "visual attention". These maps show that the model recovers spatial specificity but with semantic meaning: it activates strongly and exclusively on CNT particles while actively suppressing background regions, and imaging artifacts like TEM grid bars. This selective focus emerges naturally from self-supervised pre-training, confirming the emergent segmentation properties previously observed in DINO-based architectures8, 14, which effectively learn to "look" only at the relevant nanomaterial even without explicit segmentation training.
@@
To visualize how the MLP classifier transforms DINOv2’s foundation model features into highly discriminative representations, we applied t-SNE 43, 44, 45 dimensionality reduction to embeddings at three stages of the classification pipeline (Figure 8). The raw 3840-dimensional DINOv2 hypercolumn features (Figure 8a) already exhibit meaningful structure: the four CNT morphology classes form loosely separated clusters, with Fiber and Matrix samples occupying relatively distinct regions while Fiber and Cluster samples show greater overlap, consistent with their morphological similarity in a few images. This initial separation validates that DINOv2’s self-supervised pre-training on natural images transfers effectively to nanomaterial microscopy, capturing morphology-relevant features without task-specific fine-tuning. However, substantial inter-class overlap remains, particularly between Fiber and Cluster samples, and between Matrix and MatrixSurface samples explaining the 2% classification error in the confusion matrix. The first MLP layer (512 dimensions, Figure 8b) significantly refines this structure: class clusters become more compact and better separated, with decision boundaries beginning to form between previously overlapping regions. The most dramatic transformation occurs after the second MLP layer (128 dimensions, 8c), where all four classes form tight, well-separated manifolds with minimal overlap. This progression demonstrates that the MLP learns to amplify task-relevant dimensions of the DINOv2 feature space while suppressing irrelevant variations, effectively discovering a low-dimensional embedding optimized for CNT morphology discrimination. The small number of outlier points visible in each cluster correspond precisely to the misclassified samples in the confusion matrix, representing genuinely ambiguous borderline cases rather than random errors. Notably, even in the final 128-dimensional space, MatrixSurface and Cluster outliers tend to drift toward each other’s clusters, while Fiber and Matrix remain distinctly separated, confirming that the 2% Cluster-to-MatrixSurface confusion reflects true morphological ambiguity along a continuum rather than model failure. This visualization provides interpretable evidence that our model’s 95.5% test accuracy stems from learning a geometrically meaningful feature space where morphological similarity corresponds directly to proximity in the embedding, rather than memorizing superficial image patterns.